The current mass of the atmosphere is approximately 5.3×1018 kg. How much fossil fuel, expressed as kilograms of carbon (C), must have been burned to raise the CO2 concentration of the atmosphere from 210 ppmv to 330 ppmv? The bulk density of carbon is 2267 kg/m³. What is the volume in cubic kilometres of this much carbon?
Answer #1

Hi
Given the current mass of the atmosphere is 5.3*10^18 kg.
We may assume that atmosphere is made up of air and is at ambient conditions because all the air lies in the lower atmosphere.

Considering air as a mixture of nitrogen and oxygen we can say that atomic mass of air will be close to 29 and thus assuming the ideal gas we ca write.

Density =PM/RT
Here P is the pressure that we can take close to the atmospheric pressure and M is the molecule weight of the gas.
R is the universal gas conatnt and T is the temperature in K .

Using ambient condition we know the volume of one mole of the gas will be close to 24.4 L/mole.
So the total volume of the atmosphere =

Moles of air in atmosphere *volume per mole of atmosphere

Moles of air in atmosphere = Mass of atmosphere/molecular weight
5.3*10^21 /29 moles
1.82 *10^20 moles
Volume = 1.83*10^20 * 22.4 Litres
=4.4*10^18 m^3

Initial concentration of CO2 =210 ppmv
Final concentration of CO2 = 330 ppmv

Mass addition of CO2 = density *( Volume addition )
ppmv is nothing but 1 part volume in 10^6 part volume.
So the volume of CO2 initially =(4.4*10^12 )*210 m^3.
Volume of CO2 finally
= (4.4*10^12)*330 m^3.
Volume change of CO2.
=4.4*10^12*110 m^3
=4.84*10^14 m^3

Mass addition of CO2 = density*4.84*10^14 m^3
Density of CO2 = Density of ideal gas at the ambient conditions.

Density =PM/RT =101325*44*0.001/8.314*298
=1.79 kg/m^3
=8.66*10^14 kg
We know in the production of 44 g CO2 we need 12 gram of carbon.
So the equivalent mass of carbon required=

(8.66*12/44 )*10^14
=2.36*10^14 kg
Volume of carbon =Mass /density
=2.36*10^14/2267 m^3.
=1.04*10^11 m^3.
1km^3 = 10^9 m^3.
So the volume
=104 km^3.

Hope this helps

Thanks